### v1

@@ 6 Threats to Validity
External Validity: refers to the extent to which research findings can be generalized to broader populations or situations [145]. In this work, we analyzed Stack Overflow and GitHub repositories to understand the challenges of SWSs development. First, we used SO data to identify the challenges faced by developers of SWSs. However, this data might not encompass all the difficulties experienced by developers, as there could be subtle aspects rarely discussed. To address this limitation, we also examined prevalent issues and pull requests using GitHub, identifying common topics discussed there. Developers may also communicate their discussions through other media (e.g., mailing lists and SWSs forums). Future work considering other data sources may complement our study. Furthermore, another issue arises from the limited selection of only 11 SWSs based on their high prevalence on Stack Overflow. This selection may not represent the broader population of SWSs, as it focuses only on the most discussed systems, potentially overlooking less popular or emerging SWSs that could exhibit different characteristics or trends. The findings derived from these 11 systems may not generalize to all SWSs, especially those that are less documented or used by smaller, specialized communities. Consequently, the conclusions drawn might reflect biases inherent in the selection process, limiting the applicability of the results to the broader landscape of scientific workflow systems.
@@
7 Conclusion
This paper explores the challenges faced by developers of scientific workflow systems by analyzing relevant Stack Overflow posts and examining GitHub issues and pull requests. The findings reveal that developers encounter traditional software engineering challenges, such as data structures, operations, and dependency management, as well as unique challenges specific to SWSs, including workflow creation and task scheduling. Some SWS-related areas (e.g., workflow creation, execution, task scheduling, data structures, and operations) are gaining the highest attention from developers while being very challenging to them. We believe our findings provide valuable insights for the SWSs community by identifying areas that require focused research, particularly in addressing the most challenging aspects of SWSs development. Additionally, our results can assist SWS platform owners in gaining a better understanding of the challenges developers face and implementing strategies to address these issues effectively. Educators in the SWSs domain can also benefit from this study by incorporating the identified topics and their associated difficulties into their course design, ensuring that their courses are aligned with the real-world challenges encountered by developers.
Our study provides valuable insights for researchers and practitioners in the SWSs community, shedding light on the challenges faced by SWSs developers. However, we plan to extend our research by examining developer discussions across SWSs forums to draw more accurate and generalizable conclusions. Our future work will involve a comprehensive investigation of all SWSs forums, repositories, and posts to gain deeper insights into the various issues developers encounter and their strategies for resolving them.
Acknowledgments

### v2

@@ 6 Threats to Validity
External Validity: refers to the extent to which research findings can be generalized to broader populations or situations [164]. In this work, we analyzed Stack Overflow and GitHub repositories to understand the challenges of SWSs development. First, we used SO data to identify the challenges faced by developers of SWSs. However, this data might not encompass all the difficulties experienced by developers, as there could be subtle aspects rarely discussed. To address this limitation, we also examined prevalent issues and pull requests using GitHub, identifying common topics discussed there. Developers may also communicate their discussions through other media (e.g., mailing lists and SWSs forums). Developers working in proprietary environments, private organizations, or internal systems may encounter additional challenges that are not reflected in public discussions. Future work considering other data sources may complement our study. Furthermore, another issue arises from the limited selection of only 11 SWSs based on their high prevalence on Stack Overflow. This selection may not represent the broader population of SWSs, as it focuses only on the most discussed systems, potentially overlooking less popular or emerging SWSs that could exhibit different characteristics or trends. The findings derived from these 11 systems may not generalize to all SWSs, especially those that are less documented or used by smaller, specialized communities. Consequently, the conclusions drawn might reflect biases inherent in the selection process, limiting the applicability of the results to the broader landscape of scientific workflow systems.
Additionally, although our study covers SWSs used in multiple domains, including bioinformatics, genomics, cheminformatics, climate science, epidemiology, and machine learning, our findings are generalized across all systems. Specific domains may introduce unique challenges that are not fully captured in this study. For example, climate modeling workflows often require large-scale, distributed computing, whereas bioinformatics workflows emphasize data reproducibility and integration with specialized tools. A deeper domain-specific analysis would further strengthen the applicability of our findings.
@@
One potential threat to construct validity arises from the inclusion of Apache Airflow in our analysis. Although Airflow is a widely recognized workflow orchestration tool, it is also used in industrial and business-oriented contexts. Consequently, some posts related to Airflow in our dataset may reflect challenges associated with general-purpose workflow automation rather than scientific workflows. While we believe that Airflow remains relevant given the increasing overlap between scientific data processing practices and industrial data engineering, we acknowledge that this broader usage context may introduce minor noise into our characterization of scientific workflow development challenges.
Our study relies on publicly available discussions from Stack Overflow and GitHub, which do not explicitly indicate the research domain or expertise levels of users. As a result, while we analyzed widely discussed challenges across different Scientific Workflow Systems (SWSs), we may not capture domain-specific nuances or variations in challenges based on user proficiency. Future studies could integrate metadata analysis or user profiling to better distinguish challenges faced by specific domains and varying levels of expert users.
@@
7 Conclusion & Future Work
In this study, we conducted a comprehensive empirical analysis to uncover the challenges faced by developers of SWSs. By mining and modeling over 35,000 Stack Overflow posts and 162,000 GitHub issues and pull requests, we identified recurring topics of concern, assessed their difficulty, and analyzed the nature of developer questions across platforms. Our investigation revealed 10 dominant topics on Stack Overflow and 13 on GitHub, with challenges in workflow creation, execution, and distributed task management being especially prevalent. Fine-grained topic decomposition further exposed critical pain points, such as DAG configuration issues, Kubernetes-based deployment failures, and cross-task data handling inconsistencies. We found that ’How’-type questions dominate developer queries, underscoring a widespread need for procedural guidance. Difficulty analysis indicated that topics involving large-scale orchestration, infrastructure integration, and API migrations are particularly hard to resolve, often with prolonged resolution times or high unresolved rates.
These findings carry important implications for tool designers, documentation authors, and the broader SWS research community. They suggest the need for more robust abstractions, improved deployment support, and clearer documentation practices to reduce friction in SWS development. Our methodology also provides a replicable, lightweight approach for continuously monitoring developer pain points using public data, offering actionable insights to improve usability, reliability, and maintainability in the evolution of SWSs.
Future research will explore the use of prompt-based large language models (LLMs) to assist in automated workflow creation, debugging, and documentation, potentially reducing the technical burden on developers and end-users. To broaden our understanding of real-world challenges in SWS development and usage, we also plan to expand our analysis beyond Stack Overflow and GitHub by mining discussions from platforms such as WorkflowHub, Slack communities, and other relevant developer forums. In parallel, we aim to conduct interviews with researchers, workflow system developers, and research software engineers to identify critical pain points and design requirements. To complement these qualitative insights, we will deploy surveys targeting both novice and experienced SWS practitioners to gather quantitative data on usability barriers, perceived difficulties, and satisfaction with existing solutions. Furthermore, we will perform in-depth case studies across scientific domains such as bioinformatics, geosciences, and machine learning, examining workflow repositories and community discussions to surface domain-specific concerns.
8 Declarations
Funding
This research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), and by the industry-stream NSERC CREATE in Software Analytics Research (SOAR)
Ethical approval
Not Applicable
Informed consent
Not Applicable
Author Contributions
  • Khairul Alam led the conceptualization and design of the study, performed the data collection and preprocessing, conducted the topic modeling and empirical analysis, and drafted the manuscript.
  • Banani Roy contributed to the study design, provided methodological guidance, reviewed and refined the data collection and topic analysis procedures, and critically revised the manuscript.
  • Chanchal K. Roy supervised the overall research process, provided strategic direction for the study, contributed to the interpretation of results, refined topic analysis procedures, and helped with manuscript revision and finalization.
  • Kartik Mittal assisted in data acquisition and analysis, contributed to reviewing and validating results, and participated in reviewing.
@@ 8 Declarations
Clinical Trial Number in the manuscript
Not Applicable
Acknowledgments